# Cosmology using numerical relativity

### v1

@@ 2.2 Numerical relativity for cosmologists
A key difficulty in any numerical simulation is the resolution of different length scales, since computers have a finite precision with which they can represent numbers. For example, if we want to study small length scale effects over long timescales (relative to the light crossing time of the physical length scale), it will be challenging. This is both a feature and a bug, as we have already seen that NR is most needed where the length scales of interest are similar and therefore perturbative approaches fail. However, there will still be many cases where the length scales in the problem change significantly in space or time, and therefore the numerics can break down. In particular, cosmological spacetimes with local overdensities introduce an additional global scale (i.e. the Hubble parameter $H$) into the numerical system. This creates a challenge when our goal is to study local physics with a different scale. In many cases we expect strong backreaction to be important on small scales, and while it is possible that the dynamics can decouple from the background cosmological evolution, there are intermediate cases where both scales need to be fully resolved. In cases where we need to follow a range of scales in different regions, one can use Adaptive Mesh Refinement (AMR) techniques to gain one or two orders of magnitude, but it is even better to use clever choices of the evolution variables that scale out expansion, as in the tetrad formulation of Uggla et al. [40] and BSSN-like formulations as used by Zilhao et al. [45]. See also Daverio et al. [46] for a proposed scheme for cosmological spacetimes that eliminates the fluid components through the constraint equations.
@@ 2.3 Boundary conditions
The most common and straight-forward solution is to apply periodic boundary conditions, usually with the idea of treating the computational domain as being a typical patch of the entire larger universe, rather than being truly topologically a torus on the relevant scales, see Fig. 6. In late-time post-Big Bang Nucleosynthesis cosmological evolution this is a well-justified choice as the spacetime is observed to be statistically homogeneous and isotropic [^5] Specifically, correlation functions $\langle XY\dots\rangle$ are invariant under translation and rotation.. In the early universe the choice is less easy to justify – we simply do not know what the topology of our universe is on scales beyond the cosmological horizon. Imposing periodic boundary conditions in Cartesian coordinates restricts the topology of space to be a 3-Torus, and hence the Ricci scalar of the spatial hypersurface ${}^{(3)}R$ has to be negative somewhere or ${}^{(3)}R=0$ everywhere if there are no BHs [47]. Operationally, periodic boundary conditions also impose restrictions on the choice of initial conditions [48] – we will discuss this in Sec. 2.4 later. Finally, periodic boundary conditions impose a discrete symmetry in the size of the domain with a physical scale that is potentially evolving since it depends on the gauge. Thus one must ensure that the scale of the physics being investigated is smaller than the scale of this discrete symmetry, or accept that the results may depend on the choice of boundaries. This issue has recently been raised in the context of studies of inflation in [49], and more generally in N-body studies of cosmological evolution in [50].
@@
Alternative approaches have been taken, in which one instead matches to a particular FLRW spacetime asymptotically, see for example [51, 52, 53]. There is unfortunately no guarantee that such a choice will be consistent – imagine we choose to match to an asymptotically de Sitter region during a simulation of inflation [54], and some dynamics of the field at the interior of the domain propagate waves out to the boundary. This should disturb the boundary in some way and so rigid de Sitter becomes inconsistent and creates unphysical reflections of the waves. One may be able to fix such a case by applying de Sitter plus outgoing waves as is done for Sommerfeld boundary conditions in an asymptotically flat spacetime. However, the problem is more general – the choice of boundaries has a physical impact if the timescales of interest allow them to communicate with the physical region of interest, and therefore their choice is part of the specification of the physical situation under study.
@@
This is an area in which further work is needed, and new ideas would be welcome. Some options might be to revisit alternative boundary choices that have previously been explored in NR, such as absorbing or constraint preserving boundary conditions [55, 56, 57, 58, 59, 60], to try to compactify the domain as in recent work for gravitational wave extraction [61, 62, 63, 64], or to impose that the boundary tracks a cosmological horizon or restricts the domain to that bounded by the null lightcones emanating from the boundaries.
@@ 2.4 Initial data for cosmological simulations
In more general cases, one in principle has to solve the coupled set of constraints Eqn. (26). Both the CTT and the CTS methods that were initially developed for astrophysical simulations such as BH and NS mergers have been successfully used in cosmological NR simulations. In the former, the freely specified variables are $\{\bar{\gamma}_{ij},K,\bar{A}^{ij}_{\mathrm{TT}}\}$ together with the gauge variables $\alpha$ and $\beta^{i}$, while in the latter the freely specified variables are $\{\bar{\gamma}_{ij},\partial_{t}\gamma_{ij},K,\partial_{t}K\}$ and $\alpha$ and $\beta^{i}$ are the determined variables.
@@
The presence of fundamental matter fields plays a crucial role in many cosmological simulations. In such cases, matter is represented by one or more fundamental fields such as the scalar inflaton, which presents a special challenge in setting initial conditions. This is because unlike fluids, which can be represented by macroscopic variables such as energy density $\rho$ and pressure $p$, these variables themselves are functions of the fields and the metric. For example, for a single scalar field $\phi$ with potential $V(\phi)$, the energy density and momentum are
@@
where the source $\sigma(\psi_{0})=-2\psi_{0}^{5}\rho-\partial^{j}\partial_{j}\psi_{0}+(1/12)\psi_{0}^{5}K^{2}$. If the boundary is Dirichlet and $q(x)>0$ everywhere in the domain and its boundary, then the maximal principle implies that the only solution is $u=0$ and hence a unique solution for $\psi$ in Eqn. (32) exists. Conversely, if $q(x)<0$ somewhere, then this cannot be guaranteed and a unique solution may not exist. A standard trick to render the solution unique is to rescale $\rho\rightarrow\tilde{\rho}\psi^{-5}$, which switches the sign of the $\rho$ term in $\sigma$ and ensures that $q(x)>0$. If $\rho$ itself depends on $\psi$ and $\phi$, such a rescaling may not always allow $\phi$ to be simply reconstructed, although in Corman & East [72] a rescaling of the field itself was successfully employed for a massive potential, along with bespoke analytic solutions that ensure the momentum constraint is always satisfied as proposed by Garfinkle & Mead [67].
If rescaling is not possible, one can still attempt to numerically solve the constraints to find a solution (if it exists) – if a trial solution is close to a solution and the perturbations are weak, this method can still work. Alternatively, in Aurrekoetxea et. al. [48], it was proposed that instead of specifying $K$ and solving the Hamiltonian constraint Eqn. (32) for $\psi$ as an elliptic equation, one instead specifies $\psi$ and then solves for $K$ as an algebraic equation (up to a sign). This so-called CTTK method has been successfully employed in cosmological NR problems [73, 74, 75, 76] and was inspired by the Friedmann equations where $H\sim K$ varies with $\rho$. It also works for BBH scenarios [77, 78], but note that as a result of the spatial variation in $K$, solving the momentum constraint cannot be avoided.
@@ 2.5 Gauge choices
There are two key considerations that set the requirements for a choice of gauge in numerical simulations – firstly, and most importantly, is the stability of the evolution. The second consideration is that we would like to efficiently sample the spacetime, and want to distribute the coordinates across the region of interest in a way that captures the relevant physical scales, but does not oversample them. As discussed above, in a BSSN-type formalism, the gauge freedom is broadly encapsulated in the lapse and shift functions, $\alpha$ and $\beta^{i}$.
@@
where $\eta>0$ is a dimensionful constant of order $\sim 1/M$ that damps the shift evolution, and $B^{i}$ (not to be confused with the $B_{i}$ variable in Eqn. (7)) is an auxiliary variable. The Gamma-driver avoids slice stretching by seeking to minimise $\tilde{\Gamma}^{i}\sim-\partial_{j}\tilde{\gamma}^{ij}$, and hence preventing the formation of coordinate singularities. Using Eqn. (36) with $f(\alpha)=1/\alpha$ ($1+\log$ slicing) and Eqn. (37) combined is known as the moving puncture gauge, which allows for the stable evolution of moving BHs. This gauge permitted the successful simulations of BH mergers [88, 89], without the explicit excision of singularities that is required in the GHC formalism [^8] A common gauge used in analytic solutions of GR is the harmonic gauge, which imposes a condition on the coordinates $x^{\mu}$ such that they satisfy the equation $\Box x^{\mu}=0~{},$ (38) In the GHC formalism, the idea is to drive this wave equation with a set of source functions $H^{\mu}$ [39, 37] $\Box x^{\mu}=H^{\mu}~{},$ (39) where $H^{\mu}$ is evolved dynamically with a set of differential operators ${\cal L_{(\mu)}}$ that can be freely specified as a function of the spacetime coordinates, the metric and its derivatives, and the source functions and their derivatives. Their exact form depends on the physics being investigated and can be chosen in such a way as to achieve slicings similar to those discussed for the moving puncture gauge. Eqn. (39) and its derivatives then form a set of constraint equations, which can be substituted into the Einstein equations to form a set of well-posed equations of motion for the metric. [37, 38]. Both the standard GHC and moving-puncture gauges have been successfully applied to NR evolutions of cosmological spacetimes, including cases in which BHs form [90, 91, 92, 93, 72] [^9] Recently, an alternative shock-avoiding choice of $f(\alpha)=1+\kappa/\alpha^{2}$, first proposed by Alcubierre [94, 95], has been found to be efficient in evolving spacetimes that contain collapsing matter [96], but this has yet to be applied to cosmological evolutions..
@@
where $\langle K\rangle$ is the spatial average of $K$ over the hyperslice – the idea is to set the rate of change of $\alpha$ such $\partial_{t}{\alpha}=0$ when space is homogeneous, since $K=\langle K\rangle$. This gauge choice approximately ensures that simulation time is identified with the cosmic time coordinate rather than conformal time. A similar gauge was used by Kou et al. [98], given by
@@
Again, the idea is to prevent a runaway growth of $\alpha$ in homogeneous (and hence expanding) regions, while still avoiding the focusing of observers via the Bona-Masso condition in collapsing regions.
An alternative approach in the tetrad formulation proposed by Uggla et al. [40], requires choosing constant mean curvature slices with homogeneous (but time-dependent) $K(t)$ to foliate the spacetime. Such a choice tipically requires solving an elliptic equation for the lapse $\alpha$ at every numerical time step, since the aim is to achieve
@@
which is potentially more computationally expensive and may lead to issues of uniqueness and existence of solutions. Nevertheless, it has the value of possessing a clear physical interpretation – these are observers for which local expansion is homogeneous even if the matter distribution is not (though clearly these will not generically be geodesic observers). Other interesting recent propositions are the use of mean curvature flow, as in [99], or a reference metric approach as in [100]. The former requires implicit timestepping methods, which are less commonly used than explicit ones. However, the additional technical complexity may be worthwhile for achieving better gauge behavior in certain cases.
@@ 2.6 Interpretation of quantities in an inhomogeneous spacetime
One issue is that for any higher-ranked tensor beyond the spacetime scalar, the component values depend on the basis choice – for example for $X_{\mu}$ the physical meaning of $X_{0}$, $X_{1}$ etc depends on what basis the index $\mu$ is associated with. Even when accounting for that (or using true spacetime scalars [101, 49]), what you see on the numerical grid at some coordinate time $t$ still depends on the choice of foliation. This is particularly important when comparing the numerical results from different simulations. In principle, once a simulation is complete, one can “refoliate the spacetime” via coordinate transforms into any foliation one might desire, but in practice this is extremely difficult. It may also be the case, given the finite extent of the domain in time and space, that simulations with different foliations capture substantially different regions of the spacetime.
@@
An important distinction is between what we will refer to as gauge dependence and slicing dependence. Consider the specific example of the energy density of the ADM decomposition $\rho=n^{\mu}n^{\nu}T_{\mu\nu}$ where $n^{\mu}$ is the unit vector orthonormal to the hypersurfaces of the foliation. This is a spacetime scalar, and therefore it must be gauge independent. However, it is the energy density measured by a particular observer, specifically the one following the normal congruence $n^{\mu}$, and therefore if we change our slicing, and hence the physical trajectory that we label $n^{\mu}$, we change the physical scalar that we calculate. Therefore the value of $\rho$ that we measure depends on the slicing – it is slicing dependent. If we were able to track the original normal observer in the new gauge, and calculate the energy density associated with them, then the value would remain the same, but this is not usually done in practise. Mapping between specific observers in inhomogeneous spacetimes is not straightforward (especially if one changes the initial data), and a simulation in a different gauge will normally be associated with a different set of observers, and their associated observables.
@@
A related problem is the tendency to use average quantities across the spatial slices as some measure of the evolution. We must be careful in interpreting these since there is no possible time-like observer who can measure such quantities – only a god-like observer who can freeze time and move around the spatial slice would ever be able to measure a quantity at space-like separated points. Even if the quantity is a spacetime scalar, its integral over the surface is slicing dependent, and so the time evolution may say more about the gauge than the physics. In particular, seeing some quantity blow up can often be a sign of a coordinate singularity forming, rather than a physical instability. This observation is particularly important in the context of quantifying cosmological observations, which are made on a lightcone and not spatial hypersurfaces.
@@
In NR simulations, GWs are often extracted by projecting $\tilde{A}_{ij}$ into the TT gauge
@@
where ;0{;0}; 0 is the covariant derivative compatible with the background metric. In Fig. 7 we show their results, where they compared the extracted $\Omega_{\mathrm{GW}}$ from oscillon preheating using a rigid FLRW scheme and NR simulations with different gauges: (i) synchronous gauge (geodesic slicing): $\alpha=1$ and $\beta^{i}=0$; (ii) A combination of the 1+log slicing and Gamma driver shift conditions, Eqn. (40); (iii) Radiation gauge in Eqn. (41). This comparative approach aimed to understand the impact of NR gauge choices on the extraction of gravitational waves in a fully relativistic setting, and they concluded that the extracted GWs are highly sensitive to the choice of gauge. Synchronous gauge (geodesic slicing) was found to be an unreliable gauge condition for extracting GW spectra, as it contained redundant gauge modes that artificially enhanced the GW power spectrum by 1-2 orders of magnitude.
This example highlights the importance of carefully choosing and interpreting diagnostics from NR simulations.

### v2

@@ 2.2 Numerical relativity for cosmologists
A key difficulty in any numerical simulation is the resolution of different length scales, since computers have a finite precision with which they can represent numbers. For example, if we want to study small length scale effects over long timescales (relative to the light crossing time of the physical length scale), it will be challenging. This is both a feature and a bug, as we have already seen that NR is most needed where the length scales of interest are similar and therefore perturbative approaches fail. However, there will still be many cases where the length scales in the problem change significantly in space or time, and therefore the numerics can break down. In particular, cosmological spacetimes with local overdensities introduce an additional global scale (i.e. the Hubble parameter $H$) into the numerical system. This creates a challenge when our goal is to study local physics with a different scale. In many cases we expect strong backreaction to be important on small scales, and while it is possible that the dynamics can decouple from the background cosmological evolution, there are intermediate cases where both scales need to be fully resolved. In cases where we need to follow a range of scales in different regions, one can use Adaptive Mesh Refinement (AMR) techniques to gain one or two orders of magnitude [^8] See Radia et al. [75] for a detailed description and application of AMR techniques in numerical relativity., but it is even better to use clever choices of the evolution variables that scale out expansion, as in the tetrad formulation of Uggla et al. [70] and BSSN-like formulations as used by Zilhao et al. [76]. See also Daverio et al. [77] for a proposed scheme for cosmological spacetimes that eliminates the fluid components through the constraint equations.
A crucial step in any numerical simulation is convergence testing. These tests are often performed using three simulations with low, medium and high resolutions $\Delta_{\mathrm{LR}}>\Delta_{\mathrm{MR}}>\Delta_{\mathrm{HR}}$. The relative error of a quantity $u$ between simulations should then decrease when increasing the resolution as
$\lim_{\Delta\rightarrow 0}\frac{|u_{\mathrm{HR}}-u_{\mathrm{MR}}|}{|u_{\mathrm{MR}}-u_{\mathrm{LR}}|}\approx\frac{\Delta_{\mathrm{HR}}^{n}-\Delta_{\mathrm{MR}}^{n}}{\Delta_{\mathrm{MR}}^{n}-\Delta_{\mathrm{LR}}^{n}}\,,$ (28)
where $n$ is the order of convergence, which should match the order of the finite difference stencils used, see [58] for more details. When the exact solution of a quantity is known, two resolutions $\Delta_{\mathrm{LR}}>\Delta_{\mathrm{HR}}$ are sufficient. In NR, this allows convergence tests of the Hamiltonian and momentum constraints to be carried out using two simulations as
$\lim_{\Delta\rightarrow 0}\frac{\mathcal{H}_{\mathrm{HR}}}{\mathcal{H}_{\mathrm{LR}}}=\lim_{\Delta\rightarrow 0}\frac{\mathcal{M}_{\mathrm{HR}}}{\mathcal{M}_{\mathrm{LR}}}\approx\left(\frac{\Delta_{\mathrm{HR}}}{\Delta_{\mathrm{LR}}}\right)^{n}\,,$ (29)
where $\Delta\rightarrow 0$ indicates that this limit is expected to hold only when the resolution is sufficiently high. We emphasise here that the smallness of the constraint violation is not particularly meaningful in isolation, although one can compare it to the sum of the absolute magnitudes of the terms that appear in the constraints, as an approximate measure of the typical energy or momentum scales in the simulation (which may change over time in an evolving cosmology). The most important test is that it remains bounded in the simulation (i.e. with slower than exponential growth) and converges at the appropriate order when resolution is increased. It is common to excise regions within apparent horizons and at the boundaries (where constraint violation is inevitable but should not affect the region of interest) when assessing this convergence.
@@ 2.3 Boundary conditions
The most common and straight-forward solution is to apply periodic boundary conditions, usually with the idea of treating the computational domain as being a typical patch of the entire larger universe, rather than being truly topologically a torus on the relevant scales, see Fig. 6. In late-time post-Big Bang Nucleosynthesis cosmological evolution this is a well-justified choice as the spacetime is observed to be statistically homogeneous and isotropic [^9] Specifically, correlation functions $\langle XY\dots\rangle$ are invariant under translation and rotation.. In the early universe the choice is less easy to justify – we simply do not know what the topology of our universe is on scales beyond the cosmological horizon. Imposing periodic boundary conditions in Cartesian coordinates restricts the topology of space to be a 3-Torus, and hence the Ricci scalar of the spatial hypersurface ${}^{(3)}R$ has to be negative somewhere or ${}^{(3)}R=0$ everywhere if there are no BHs [78]. Operationally, periodic boundary conditions also impose restrictions on the choice of initial conditions [79] – we will discuss this in Sec. 2.4 later. Finally, periodic boundary conditions impose a discrete symmetry in the size of the domain with a physical scale that is potentially evolving since it depends on the gauge. Thus one must ensure that the scale of the physics being investigated is smaller than the scale of this discrete symmetry, or accept that the results may depend on the choice of boundaries. This issue has recently been raised in the context of studies of inflation in [80], and more generally in N-body studies of cosmological evolution in [81, 82].
@@
Alternative approaches have been taken, in which one instead matches to a particular FLRW spacetime asymptotically, see for example [83, 84, 85, 86]. If compactified coordinates are used as in [85, 86] it is possible to consistently impose the boundary conditions at asymptotic infinity, but since this imposes the homogeneous solution at sufficiently large distances, it removes any “global” backreaction from the inhomogeneities. In other codes boundary conditions are imposed at large but finite distances from the system of interest. In such a case, there is no guarantee that such a choice will be consistent – imagine we choose to match to an asymptotically de Sitter region during a simulation of inflation [87], and some dynamics of the field at the interior of the domain propagate waves out to the boundary. This should disturb the boundary in some way and so rigid de Sitter becomes inconsistent and creates unphysical reflections of the waves. One may be able to fix such a case by applying de Sitter plus outgoing waves as is done for Sommerfeld boundary conditions in an asymptotically flat spacetime. However, the problem is more general – the choice of boundaries has a physical impact if the timescales of interest allow them to communicate with the physical region of interest, and therefore their choice is part of the specification of the physical situation under study.
@@
This is an area in which further work is needed, and new ideas would be welcome. Some options might be to revisit alternative boundary choices that have previously been explored in NR, such as absorbing or constraint preserving boundary conditions [88, 89, 90, 91, 92, 93, 94, 95, 96], to try to compactify the domain as in Corman et al. [85] and recent work for gravitational wave extraction [97, 98, 99, 100], or to impose that the boundary tracks a cosmological horizon or restricts the domain to that bounded by the null lightcones emanating from the boundaries.
@@ 2.4 Initial data for cosmological simulations
In more general cases, one in principle has to solve the coupled set of constraints Eqn. (26). Both the CTT and the CTS methods [101] (or now more commonly the extended CTS method (XCTS) [102, 103]) that were initially developed for astrophysical simulations such as BH and NS mergers have been successfully used in cosmological NR simulations. In the former, the freely specified variables are $\{\bar{\gamma}_{ij},K,\bar{A}^{ij}_{\mathrm{TT}}\}$ together with the gauge variables $\alpha$ and $\beta^{i}$, while in the XCTS case the freely specified variables are $\{\bar{\gamma}_{ij},\partial_{t}\gamma_{ij},K,\partial_{t}K\}$ and $\alpha$ and $\beta^{i}$ are the determined variables.
@@
The presence of fundamental matter fields plays a crucial role in many cosmological simulations. In such cases, matter is represented by one or more fundamental fields such as the scalar inflaton, which presents a special challenge in setting initial conditions. This is because unlike fluids, which can be represented by macroscopic variables such as energy density $\rho$ and pressure $p$, these variables themselves are functions of the fields and their gradients. For example, for a single scalar field $\phi$ with potential $V(\phi)$, the energy density and momentum are
@@
where the source $\sigma(\psi_{0})=-2\psi_{0}^{5}\rho-\partial^{j}\partial_{j}\psi_{0}+(1/12)\psi_{0}^{5}K^{2}$. If the boundary is Dirichlet and $q(x)>0$ everywhere in the domain and its boundary, then the maximal principle implies that the only solution is $u=0$ and hence a unique solution for $\psi$ in Eqn. (34) exists. Conversely, if $q(x)<0$ somewhere, then this cannot be guaranteed and a unique solution may not exist. A standard trick to render the solution unique is to rescale $\rho\rightarrow\tilde{\rho}\psi^{-5}$, which switches the sign of the $\rho$ term in $\sigma$ and ensures that $q(x)>0$. If $\rho$ itself depends on $\psi$ and $\phi$, such a rescaling may not always allow $\phi$ to be simply reconstructed, although in Corman & East [109], which developed the CTS method of [110], a method of rescaling the momentum that preserved the ratio of the gradient and kinetic energy in the solution was successfully employed for a massive potential, along with bespoke analytic solutions that ensure the momentum constraint is always satisfied as proposed by Garfinkle & Mead [104]. Corman & East [109] show that for the specific choices they made unique solutions exist, although in more general cases the uniqueness of the Hamiltonian constraint is potentially problematic.
If rescaling is not possible, one can still attempt to numerically solve the constraints to find a solution (if it exists) – if a trial solution is close to a solution and the perturbations are weak, this method can still work. Alternatively, in Aurrekoetxea et. al. [79], it was proposed that instead of specifying $K$ and solving the Hamiltonian constraint Eqn. (34) for $\psi$ as an elliptic equation, one instead specifies $\psi$ and then solves for $K$ as an algebraic equation (up to a sign). This so-called CTTK method has been successfully employed in cosmological NR problems [111, 112, 113, 114] and was inspired by the Friedmann equations where $H\sim K$ varies with $\rho$. It also works for BBH scenarios [115, 116, 117], but note that as a result of the spatial variation in $K$, solving the momentum constraint cannot be avoided.
@@ 2.5 Gauge choices
There are two key considerations that set the requirements for a choice of gauge in numerical simulations – firstly, and most importantly, is the stability of the evolution. The second consideration is that we would like to efficiently sample the spacetime, and want to distribute the coordinates across the region of interest in a way that captures the relevant physical scales, but does not oversample them. As discussed above, in an ADM-based formalism, the gauge freedom is broadly encapsulated in the lapse and shift functions, $\alpha$ and $\beta^{i}$.
@@
where $\eta>0$ is a dimensionful constant (which in BH spacetimes is of order $\sim 1/M_{\mathrm{ADM}}$ but in cosmological spacetimes would be roughly set by the total mass of any matter collapsing into overdensities) that damps the shift evolution, and $B^{i}$ (not to be confused with the $B_{i}$ variable in Eqn. (7)) is an auxiliary variable. The Gamma-driver avoids slice stretching by seeking to minimise $\tilde{\Gamma}^{i}\sim-\partial_{j}\tilde{\gamma}^{ij}$, and hence preventing the formation of coordinate singularities. Using Eqn. (38) with $f(\alpha)=1/\alpha$ ($1+\log$ slicing) and Eqn. (39) combined is known as the moving puncture gauge, which allows for the stable evolution of moving BHs. This gauge permitted the successful simulations of BH mergers [127, 128], without the explicit excision of singularities that is required in the GHC formalism [^12] A common gauge used in analytic solutions of GR is the harmonic gauge, which imposes a condition on the coordinates $x^{\mu}$ such that they satisfy the equation $\Box x^{\mu}=0~{},$ (40) In the GHC formalism, the idea is to drive this wave equation with a set of source functions $H^{\mu}$ [69, 129] $\Box x^{\mu}=H^{\mu}~{},$ (41) where $H^{\mu}$ is evolved dynamically with a set of differential operators ${\cal L_{(\mu)}}$ that can be freely specified as a function of the spacetime coordinates, the metric and its derivatives, and the source functions and their derivatives. Their exact form depends on the physics being investigated and can be chosen in such a way as to achieve slicings similar to those discussed for the moving puncture gauge. Eqn. (41) and its derivatives then form a set of constraint equations, which can be substituted into the Einstein equations to form a set of well-posed equations of motion for the metric. [129, 68]. Both the standard GHC and moving-puncture gauges have been successfully applied to NR evolutions of cosmological spacetimes, including cases in which BHs form [130, 131, 132, 133, 85, 109] [^13] Recently, an alternative shock-avoiding choice of $f(\alpha)=1+\kappa/\alpha^{2}$, first proposed by Alcubierre [134, 135], has been found to be efficient in evolving spacetimes that contain collapsing matter [136], but this has yet to be applied to cosmological evolutions..
@@
where $\langle K\rangle$ is the spatial average of $K$ over the hyperslice – the idea is to set the rate of change of $\alpha$ such $\partial_{t}{\alpha}=0$ when space is homogeneous, since $K=\langle K\rangle$. This gauge choice approximately ensures that simulation time is identified with the cosmic time coordinate rather than conformal time. In GHC, something similar can be achieved by setting $H^{t}=K_{0}$, as done by East et al. [86] and Corman et al. [138]. A similar gauge was used by Kou et al. [139], given by
@@
Again, the idea is to prevent a runaway growth of $\alpha$ in homogeneous (and hence expanding) regions, while still avoiding the focusing of observers via the Bona-Masso condition in collapsing regions. Another approach to avoid violating the CFL condition due to the growing lapse, used in East et al. and Corman et al. [130, 85, 109], is adapting the timestepping by dividing it by the maximum of the lapse over the numerical domain.
An alternative approach in the tetrad formulation proposed by Uggla et al. [70], requires choosing constant mean curvature slices with homogeneous (but time-dependent) $K(t)$ to foliate the spacetime. Such a choice typically requires solving an elliptic equation for the lapse $\alpha$ at every numerical time step, since the aim is to achieve
@@
which is potentially more computationally expensive and may lead to issues of uniqueness and existence of solutions, or gauge instabilities when evolving black hole spacetimes. Nevertheless, it has the value of possessing a clear physical interpretation – these are observers for which local expansion is homogeneous even if the matter distribution is not (though clearly these will not generically be geodesic observers). Other interesting recent propositions are the use of mean curvature flow, as in [140], or a reference metric approach as in [141]. The former requires implicit timestepping methods, which are less commonly used than explicit ones. However, the additional technical complexity may be worthwhile for achieving better gauge behavior in certain cases.
@@ 2.6 Interpretation of quantities in an inhomogeneous spacetime
One issue is that for any higher-ranked tensor beyond the spacetime scalar, the component values depend on the basis choice – for example for $X_{\mu}$ the physical meaning of $X_{0}$, $X_{1}$ etc depends on what basis the index $\mu$ is associated with. Even when accounting for that (or using true spacetime scalars [142, 80, 143]), what you see on the numerical grid at some coordinate time $t$ still depends on the choice of foliation. This is particularly important when comparing the numerical results from different simulations. In principle, once a simulation is complete, one can “refoliate the spacetime” via coordinate transforms into any foliation one might desire, but in practice this is extremely difficult. It may also be the case, given the finite extent of the domain in time and space, that simulations with different foliations capture substantially different regions of the spacetime.
@@
An important distinction is between what we will refer to as gauge dependence and slicing dependence. Consider the specific example of the energy density of the ADM decomposition $\rho=n^{\mu}n^{\nu}T_{\mu\nu}$ where $n^{\mu}$ is the unit vector orthonormal to the hypersurfaces of the foliation. This is a spacetime scalar, and therefore it must be gauge independent. However, it is the energy density measured by a particular observer, specifically the one following the normal congruence $n^{\mu}$, and therefore if we change our slicing, and hence the physical trajectory that we label $n^{\mu}$, we change the physical scalar that we calculate. Therefore the value of $\rho$ that we measure depends on the slicing – it is slicing dependent. If we were able to track the original normal observer in the new gauge, and calculate the energy density associated with them, then the value would remain the same, but this is not usually done in practise. Mapping between specific observers in inhomogeneous spacetimes is not straightforward (especially if one changes the initial data), and a simulation in a different gauge will normally be associated with a different set of observers, and their associated observables. A further complication is that what we are most interested in physically is $\rho_{R}=T_{\mu\nu}u^{\mu}u^{\nu}$, the density of the fluid in its rest frame, with matter 4-velocity $u^{\mu}$, since this corresponds to the gauge-invariant density perturbations on large scales [144, 145, 146, 147]. However, since $u^{\mu}$ in general may have vorticity, it then may not be hypersurface orthogonal. Therefore a slicing corresponding to observers in the rest frame of the fluid (i.e. $n^{\mu}=u^{\mu}$) may not exist.
@@
A related problem is the tendency to use average quantities across the spatial slices as some measure of the evolution. We must be careful in interpreting these since there is no possible time-like observer who can measure such quantities – only a god-like observer who can freeze time and move around the spatial slice would ever be able to measure a quantity at space-like separated points. Even if the quantity is a spacetime scalar, its integral over the surface is slicing dependent, and so the time evolution may say more about the gauge than the physics. In particular, seeing some quantity blow up can often be a sign of a coordinate singularity forming, rather than a physical instability. This observation is particularly important in the context of quantifying cosmological observations, which are made on a lightcone and not spatial hypersurfaces. “Global” measures such as the spacetime’s Petrov classification can also be useful to understand the broad structure of the solutions, and this has been investigated by Munoz & Bruni [143, 148], although it is again hard to map them to observables.
@@
In NR simulations, GWs are often extracted by projecting $\tilde{A}_{ij}$ (defined in Eqns. (3132)) into the TT gauge
@@
where ;0{;0}; 0 is the covariant derivative compatible with the background metric. In Fig. 7 we show their results, where they compared the extracted $\Omega_{\mathrm{GW}}$ from oscillon preheating using a rigid FLRW scheme and NR simulations with different gauges: (i) synchronous gauge (geodesic slicing): $\alpha=1$ and $\beta^{i}=0$; (ii) A combination of the 1+log slicing and Gamma driver shift conditions, Eqn. (42); (iii) Radiation gauge in Eqn. (43). This comparative approach aimed to understand the impact of NR gauge choices on the extraction of gravitational waves in a fully relativistic setting, and they concluded that the extracted GWs are highly sensitive to the choice of gauge. Synchronous gauge (geodesic slicing) was found to be an unreliable gauge condition for extracting GW spectra, as it contained redundant gauge modes that artificially enhanced the GW power spectrum by 1-2 orders of magnitude. This example highlights the importance of carefully choosing and interpreting diagnostics from NR simulations. In particular, one needs to think carefully about measuring physical observables for GWs. Ideally, for example, the simulation would evolve to an approximately FLRW state at late times, so that we could then extract the locally averaged gravitational wave energy density measured by late time comoving observers. Measurement of GWs in the non-linear part of the evolution is inherently affected by gauge issues, since there is no natural “background plus perturbation” split.